A short run cost function for an entrepreneur is q3- 8q2 + 30q + 60. Determine the price at which the entrepreneur cases production in an ideal market. Also, derive the supply function.


1
Expert's answer
2021-06-07T14:41:29-0400
"C(q)=q^3-8q^2+30q+60"


"MC=3q^2-16q+30""MR=MC=>p=3q^2-16q+30""p=3(q^2-2(\dfrac{8}{3})q+\dfrac{64}{9})-\dfrac{64}{3}+30""\dfrac{p}{3}=(q-\dfrac{8}{3})^2+\dfrac{26}{9}""(q-\dfrac{8}{3})^2=\dfrac{p}{3}-\dfrac{26}{9}""q-\dfrac{8}{3}=\sqrt{\dfrac{p}{3}-\dfrac{26}{9}}"

Supply function


"q=\sqrt{\dfrac{p}{3}-\dfrac{26}{9}}+\dfrac{8}{3}"